# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v3

@@ 3.1 Content creation
Content creation is a fundamental activity of artists and designers and the term ‘AI art’ refers to artforms created with the assistance of an AI algorithms or entirely by an AI system. This can refer to various digital forms including images, texts, audio, and videos. The roots of AI art can be traced back to the 20th century, exemplified by AARON, a computer program initiated in 1972 to autonomously produce paintings and drawings [256]. The practicality of AI art has been enhanced with advancements in deep learning, particularly GANs from 2014 and, more recently, transformers, DMs and INRs.
@@ 3.1.1 Text generation, script and journalism
In the era of LLMs, AI writing tools have been widely used to assist various writing tasks, including generation written articles, blog posts, essays, and reports. These tools go beyond mere grammar and spelling checks; they boast advancements enabling them to analyze the style and tone of written material, adding images, videos and tables, offering suggestions to enhance clarity, coherence, and overall readability [257]. Moreover, AI tools extend their utility beyond content generation by automating tasks like keyword generation, meta tags, and descriptions, thereby increasing search rankings using search engine optimization (SEO). Additionally, they support the process of publishing across multiple online platforms. Transformers have been used to generate image captions by combining information from the images with a word prefix or questions [44].
@@
As discussed earlier, chatbots are now powered by LLMs, effectively simulating human conversation. These fundamental LLMs are specialized for specific tasks. For instance journalist AI and blog AI writers [^12] For example, see https://tryjournalist.com/ generate content with layouts suitable for print or online publication. Additionally, AI tools exist that are designed to detect AI-generated content (e.g., for checking for copyright), AI-writing styles, content originality and to ensure the naturalness and flow of articles. Undoubtedly, generative AI is reshaping the way artists and journalists operate. For an in-depth exploration of the impact and implications of these technological advancements on news organizations, refer to the survey conducted by Beckett et al. [261].
Generating text and scripts automatically can also be done through image and video inputs without text prompts (e.g., image captioning [262]) and with text prompts. These approaches are referred to as Vision Language Models (VLMs): multimodal models that learn from images and text. The most common and prominent models often consist of an image encoder, an embedding projector to align image and text representation, often via a dense neural network, and a text decoder stacked in this order. The most well-known technique is Contrastive Language-Image Pre-training (CLIP) [263]. More recent work in [45] scales up the vision vocabulary by incorporating new image features into the existing CLIP model, resulting in improved content understanding. A comprehensive survey of VLMs for vision tasks can be found in [264].
@@ 3.1.2 Audio and music generation
Similar to language models, AI-based music generation has rapidly advanced due to unsupervised learning on large datasets and the use of transformers (see Section 2.2). Examples of such systems include MuseNet [^13] https://openai.com/research/musenet, Magenta Studio [^14] https://magenta.tensorflow.org/studio, and Musicfy [^15] https://musicfy.lol/. These tools assist in music composition by learning complex musical patterns, predicting the next word or music note in a sequence, and mixing specified instruments. Moreover, AI tools can convert one type of sound into another, such as from whistling to violin or from flute to saxophone [^16] See an example by Ummet Ozcan at https://www.youtube.com/watch?v=lI1LCfTx2lI. This capability is invaluable for artists who may not be proficient in playing all the instruments they wish to incorporate, saving both time and costs. In 2024, Suno has released a model capable of producing radio-quality music that can be created in 2 minutes [^17] https://www.suno.ai/blog/v3. Later, Udio  [^18] https://www.udio.com/, was launched. This offers a prompt to create lyrics and music with a maximum duration of 90 seconds, and also appears to have, at least some, awareness of copyright.
AI voice software changes vocalizations from one person to another, for example enabling users to train the model to convert other people’s voices into their own, e.g. lalals [^19] https://lalals.com/, Kits [^20] https://www.kits.ai/, Media.io [^21] https://www.media.io/online-voice-changer.html, etc. Certain software, such as Voice.ai [^22] https://voice.ai/, even offers real-time voice changing capabilities. The technologies behind this uses a transformer to learn voice features and patterns in mel-spectrogram form. For example, the framework proposed in [49] uses a DM-based method with a transformer backbone to turn text input into a mel-spectrogram using the vector quantized variational autoencoder (VQ-VAE) [265]. Next, this mel-spectrogram is transformed into a sound wave. Unlike a regular spectrogram, the mel-spectrogram is based on the mel-frequency scale, which offers higher resolution for lower frequencies. Voice style transfer often uses zero-shot learning (a model is trained to recognize classes or categories that it has never encountered during training) [47] or few-shot learning (a model trained with only one or a few examples per class) [266]. Stable Audio Open [50] introduces a text-conditioned generative model for non-speech audio, trained on Creative Commons licensed data, capable of producing state-of-the-art 44.1kHz stereo audio.
@@ 3.1.4 Video generation and animation
Viggle [^41] https://viggle.ai/ claims to be the first video-3D foundation model embodying an actual understanding of physics. It combines a character and a text prompt about motion to generate character animation. Available AI tools for 3D on the market include DeepMotion [^42] https://www.deepmotion.com/ that offers text-to-3D post animation and video-to-3D post animation, shown in Fig. 4 (left). The later function can track multiple people from real video and generates replicated characters with the same motions.
@@ 3.2.1 Text categorization
Applications of text categorization include detecting spam emails, automating customer support, monitoring social media for harmful content, etc. At its core, text categorization involves assigning predefined labels to text documents, which can be anything from a tweet to a lengthy article. LLMs are particularly well-suited for this task due to their ability to comprehend complex and nuanced language. One of the main advantages of using LLMs in text categorization is their transfer learning capability. Models can be pre-trained on a large amount of text and then fine-tuned on a smaller, task-specific dataset, with or without further post-processing technique. For example, CARP [83] applies kNN to integrate diagnostic reasoning process for final decision. ChatGraph, proposed by Shi et al. [84], utilizes ChatGPT to refine text documents. It uses a knowledge graph, extracted using another specific defined prompt, and finally, a linear model is trained on the text graph for classification. Multiple learners are also used to enhance the performances [85, 86].
@@ 3.2.2 Advertisements and film analysis
Current advances in generative AI would greatly benefit sentiment analysis, also known as opinion mining, where opinions are gather from social media, articles, customer feedback, and corporate communication and are analysed to understand emotion of the owners. This is a potential tool for filmmakers and studios, enabling the creation of effective and targeted marketing campaigns. By analyzing viewer emotions and opinions, AI can provide valuable insights into audience preferences, aiding in the optimization of film marketing strategies. Sentiment analysis with modern generative AI produce more accurate results. Technically, LLMs learn complex patterns and relationships in text data for sentiment classification [87, 88]. SiEBERT [89] provides pre-trained model with open-source scripts to be fine-tuned to further improve accuracy for novel applications. Cinema Multiverse Lounge [271], a multi-agent conversational system, allows users to interact with LLM-driven agents, each embodying a distinct film-related target users.
@@ 3.2.3 Content retrieval and recommendation services
Generative retrieval (GR) was pioneered by Metzler et al. [90]. Unlike traditional retrieval, which adheres to the “index-retrieve-then-rank” paradigm, the GR paradigm employs a single model to obtain results from query input. The model generally involve deep-learning based transformers, generating output token-by-token. More recent work in [95] introduces learning-to-rank training to enhance the performance system up to 30%. GR has several advantages including substituting the bulky external index with an internal index (i.e., model parameters), significantly reducing memory usage, and enabling optimization during end-to-end model training towards a universal objective for information retrieval tasks. Conversational question answering techniques have been integrated to enhance the document retrieval [94].
@@ 3.3.1 Enhancement
In low-light conditions, scenes often exhibit low contrast, leading to focusing difficulties or the need for long exposures, which can result in blurred images and videos. To address this, LEDNet [279] has introduced a synthetic dataset for such scenarios and incorporated a learnable non-linear activation function within the network to enhance feature intensities. Meanwhile, SNR-Aware [98] estimates spatial-varying Signal-to-Noise Ratio (SNR) maps and proposes local and global learning branches using ResNet and transformer architectures, respectively. NeRCo [108] address low-light problem with INR, which unifies the diverse degradation factors of real-world scenes with a controllable fitting function. Diffusion models (DMs) have also become popular choices for low-light image enhancement [104, 105, 106]. Diff-Retinex [105] formulates the low-light image enhancement problem into Retinex decomposition, and employs multi-path generative diffusion networks to reconstruct the normal-light Retinex probability distribution. A recent state-of-the-art approach presented in [106] decomposes images into high and low frequencies using wavelet transform. High frequencies are enhanced using a transformer-based pipeline, while the low frequencies undergo a diffusion process. This method achieves nearly 2.8dB improvement over the state-of-the-art transformer-based approach, e.g. LLFormer [100], and significantly better than INR-based method, NeRCo [108], on a real low-light image benchmarking dataset. The technique has been extended for video enhancement in [107]. The output of the enhancement typically depends on user preferences. This has been viewed as a one-to-many inverse problem, with attempts to solve it using Bayesian approaches. For example, a Bayesian Enhancement Model (BEM) [280] incorporates Bayesian Neural Networks (BNNs) to capture data uncertainty and produce diverse outputs. The method can be used with Transformers or Mamba as the architecture backbone.
@@ 3.3.4 Restoration
In our previous review paper [8], we categorized the work on restoration into several different types of distortions, including deblurring, denoising, dehazing, and mitigating atmospheric turbulence. Recent work however uses a unified network architecture to address these as inverse problems $y=hx+n$, where $x$ and $y$ are the ideal and observed data, respectively. $h$ is a degradation function, such as blur, and $n$ is additive noise. Often the super-resolution task is also considered as an inverse problem, meaning $h$ includes downsampling process. Note that although designed as a single network, the model is trained with each distorted dataset separately.
@@
The pioneering transformer-based method for image restoration, SwinIR [115], employs several concatenated Swin Transformer blocks [16]. SwinIR surpasses state-of-the-art CNN-based methods proposed up to the year 2021 in super-resolution and denoising tasks. The model is smaller and reconstructs fine details more effectively. Other two popular approaches that emerged in the same timeframe are Uformer [130] and Restormer [131]. Both incorporate Transformer blocks into hierarchical encoder-decoder networks, employing skip connections similar to those in U-Net. Their objective was to restore noisy images, sharpen blurry images, and remove rain. The networks focused on predicting the residual $R$ and obtaining the restored image $\hat{x}$ through $\hat{x}=y+R$. While their performance is very similar, Restormer has half the parameters of Uformer. More recent, GRL by Li et al. [119] exploits a hierarchy of features in a global, regional, and local range using different ways to compute self-attentions as an image often show similarity within itself in different scales and areas. GRL outperforms SwinIR and Restormer. Additionally, Fei et al. introduced the Generative Diffusion Prior [128] for unsupervised learning, aiming to model posterior distributions for image restoration and enhancement. VmambaIR [143] incorporates Mamba blocks into the U-Net architecture, achieving superior performance compared to SwinIR and Restormer in both visual quality and model size.
@@
For video restoration, the general framework comprises frame alignment, feature fusion and reconstruction. The process could be similar to image restoration but input multiple frames and run through the sequences in sliding window manner to exploit temporal information of a number of consecutive frames. Recently, Video Restoration Transformer (VRT) [103] and its improved version with recurrent process (RVRT) [99], have emerged as the state of the arts for video super-resolution, deblurring, denoising, and frame interpolation. This method introduces temporal reciprocal self-attention in the transformer architecture and parallel warping using MLP. These innovations enable parallel computation and outperform the previous state-of-the-art methods by up to 2.16dB on benchmark datasets. FMA-Net [102] proposed multi-attention for joint Video super-resolution and deblurring, achieving fast runtime with nearly 40% improvement over RVRT, and the restored quality was reported better by up to 3%.
@@
iii) Dehazing: Vision transformers for single image dehazing was proposed in DehazeFormer [135]. Similar to SUNet, it is UNet-like architecture, but introduced Rescale Layer Normalization for better suit on improving contrast. The Fast Fourier Transform (FFT) has been employed in [140] due to the phase spectrum conveying more structural detail than the amplitude spectrum and demonstrating greater robustness to contrast distortion and noise. Then cross-attention between the RGB and YCbCr color spaces is applied. This approach achieves nearly 5 dB higher PSNR than DehazeFormer on a real-world smoke dataset. For video dehazing, Xu et al. [136] introduced a recurrent multi-range scene radiance recovery module with the space-time deformable attention. They also employs physics prior to inform haze attenuation. This method outperforms DehazeFormer by approximately 1dB.
iv) Mitigating atmospheric turbulence: Similar to dehazing, physics-inspired models have been widely developed to remove turbulence distortion [146, 148], while complex-valued CNNs have been proposed to exploit phase information [297]. There was also an attempt to use instance normalization (INR) to address this issue, providing solutions for tile and blur correction [148]. However, diffusion models have shown superior performance on single-image restoration tasks [145], while transformer-based methods remained the state-of-the-art for video restoration [138, 139]. More recently, Mamba-based architectures have demonstrated their effectiveness in both visual quality and model efficiency [298]. A recent review can be found in [299].

### v4

@@ 3.1 Content creation
Content creation is a fundamental activity of artists and designers and the term ‘AI art’ refers to artforms created with the assistance of an AI algorithm or entirely by an AI system. This can refer to various digital forms including images, texts, audio, and videos. The roots of AI art can be traced back to the 20th century, exemplified by AARON, a computer program initiated in 1972 to autonomously produce paintings and drawings [256]. The practicality of AI art has been enhanced with advancements in deep learning, particularly GANs from 2014 and, more recently, transformers, DMs and INRs.
@@ 3.1.1 Text generation, script and journalism
In the era of LLMs, AI writing tools have been widely used to assist various writing tasks, including generation of written articles, blog posts, essays, and reports. These tools go beyond mere grammar and spelling checks; they boast advancements enabling them to analyze the style and tone of written material, adding images, videos and tables, offering suggestions to enhance clarity, coherence, and overall readability [257]. Moreover, AI tools extend their utility beyond content generation by automating tasks like keyword generation, meta tags, and descriptions, thereby increasing search rankings using search engine optimization (SEO). Additionally, they support the process of publishing across multiple online platforms. Transformers have been used to generate image captions by combining information from the images with a word prefix or questions [44].
@@
As discussed earlier, chatbots are now powered by LLMs, effectively simulating human conversation. These fundamental LLMs are specialized for specific tasks. For instance, journalist AI and blog AI writers [^12] For example, see https://tryjournalist.com/ generate content with layouts suitable for print or online publication. Additionally, AI tools exist that are designed to detect AI-generated content (e.g., for checking for copyright), AI-writing styles, content originality, and to ensure the naturalness and flow of articles. Undoubtedly, generative AI is reshaping the way artists and journalists operate. For an in-depth exploration of the impact and implications of these technological advancements on news organizations, refer to the survey conducted by Beckett et al. [261].
Generating text and scripts automatically can also be done through image and video inputs without text prompts (e.g., image captioning [262]) and with text prompts. These approaches are referred to as Vision Language Models (VLMs): multimodal models that learn from images and text. The most common and prominent models often consist of an image encoder, an embedding projector to align image and text representations, often via a dense neural network, and a text decoder stacked in this order. The most well-known technique is Contrastive Language-Image Pre-training (CLIP) [263]. More recent work in [45] scales up the vision vocabulary by incorporating new image features into the existing CLIP model, resulting in improved content understanding. A comprehensive survey of VLMs for vision tasks can be found in [264].
@@ 3.1.2 Audio and music generation
Similar to language models, AI-based music generation has rapidly advanced due to unsupervised learning on large datasets and the use of transformers (see Section 2.2). Examples of such systems include MuseNet [^13] https://openai.com/research/musenet, Magenta Studio [^14] https://magenta.tensorflow.org/studio, and Musicfy [^15] https://musicfy.lol/. These tools assist in music composition by learning complex musical patterns, predicting the next word or music note in a sequence, and mixing specified instruments. Moreover, AI tools can convert one type of sound into another, such as from whistling to violin or from flute to saxophone [^16] See an example by Ummet Ozcan at https://www.youtube.com/watch?v=lI1LCfTx2lI. This capability is invaluable for artists who may not be proficient in playing all the instruments they wish to incorporate, saving both time and costs. In 2024, Suno has released a model capable of producing radio-quality music that can be created in 2 minutes [^17] https://www.suno.ai/blog/v3. Later, Udio [^18] https://www.udio.com/ was launched. This offers a prompt to create lyrics and music with a maximum duration of 90 seconds, and also appears to have, at least some, awareness of copyright.
AI voice software changes vocalizations from one person to another, for example enabling users to train the model to convert other people’s voices into their own, e.g., lalals [^19] https://lalals.com/, Kits [^20] https://www.kits.ai/, Media.io [^21] https://www.media.io/online-voice-changer.html, etc. Certain software, such as Voice.ai [^22] https://voice.ai/, even offers real-time voice changing capabilities. The technologies behind this use a transformer to learn voice features and patterns in mel-spectrogram form. For example, the framework proposed in [49] uses a DM-based method with a transformer backbone to turn text input into a mel-spectrogram using the vector quantized variational autoencoder (VQ-VAE) [265]. Next, this mel-spectrogram is transformed into a sound wave. Unlike a regular spectrogram, the mel-spectrogram is based on the mel-frequency scale, which offers higher resolution for lower frequencies. Voice style transfer often uses zero-shot learning (a model is trained to recognize classes or categories that it has never encountered during training) [47] or few-shot learning (a model trained with only one or a few examples per class) [266]. Stable Audio Open [50] introduces a text-conditioned generative model for non-speech audio, trained on Creative Commons licensed data, capable of producing state-of-the-art 44.1kHz stereo audio.
@@ 3.1.4 Video generation and animation
Viggle [^41] https://viggle.ai/ claims to be the first video-3D foundation model embodying an actual understanding of physics. It combines a character and a text prompt about motion to generate character animation. Available AI tools for 3D on the market include DeepMotion [^42] https://www.deepmotion.com/ that offers text-to-3D post animation and video-to-3D post animation, shown in Fig. 4 (left). The later function can track multiple people from real video and generate replicated characters with the same motions.
@@ 3.2.1 Text categorization
Applications of text categorization include detecting spam emails, automating customer support, monitoring social media for harmful content, etc. At its core, text categorization involves assigning predefined labels to text documents, which can be anything from a tweet to a lengthy article. LLMs are particularly well-suited for this task due to their ability to comprehend complex and nuanced language. One of the main advantages of using LLMs in text categorization is their transfer learning capability. Models can be pre-trained on a large amount of text and then fine-tuned on a smaller, task-specific dataset, with or without further post-processing techniques. For example, CARP [83] applies kNN to integrate a diagnostic reasoning process for final decision. ChatGraph, proposed by Shi et al. [84], utilizes ChatGPT to refine text documents. It uses a knowledge graph, extracted using another specifically defined prompt, and finally, a linear model is trained on the text graph for classification. Multiple learners are also used to enhance the performance [85, 86].
@@ 3.2.2 Advertisements and film analysis
Current advances in generative AI would greatly benefit sentiment analysis, also known as opinion mining, where opinions are gathered from social media, articles, customer feedback, and corporate communication and are analyzed to understand the emotion of the owners. This is a potential tool for filmmakers and studios, enabling the creation of effective and targeted marketing campaigns. By analyzing viewer emotions and opinions, AI can provide valuable insights into audience preferences, aiding in the optimization of film marketing strategies. Sentiment analysis with modern generative AI produces more accurate results. Technically, LLMs learn complex patterns and relationships in text data for sentiment classification [87, 88]. SiEBERT [89] provides a pre-trained model with open-source scripts to be fine-tuned to further improve accuracy for novel applications. Cinema Multiverse Lounge [271], a multi-agent conversational system, allows users to interact with LLM-driven agents, each embodying a distinct film-related target user.
@@ 3.2.3 Content retrieval and recommendation services
Generative retrieval (GR) was pioneered by Metzler et al. [90]. Unlike traditional retrieval, which adheres to the “index-retrieve-then-rank” paradigm, the GR paradigm employs a single model to obtain results from query input. The model generally involves deep-learning based transformers, generating output token-by-token. More recent work in [95] introduces learning-to-rank training to enhance the performance system up to 30%. GR has several advantages including substituting the bulky external index with an internal index (i.e., model parameters), significantly reducing memory usage, and enabling optimization during end-to-end model training towards a universal objective for information retrieval tasks. Conversational question answering techniques have been integrated to enhance the document retrieval [94].
@@ 3.3.1 Enhancement
In low-light conditions, scenes often exhibit low contrast, leading to focusing difficulties or the need for long exposures, which can result in blurred images and videos. To address this, LEDNet [279] has introduced a synthetic dataset for such scenarios and incorporated a learnable non-linear activation function within the network to enhance feature intensities. Meanwhile, SNR-Aware [98] estimates spatial-varying Signal-to-Noise Ratio (SNR) maps and proposes local and global learning branches using ResNet and transformer architectures, respectively. NeRCo [108] addresses the low-light problem with INR, which unifies the diverse degradation factors of real-world scenes with a controllable fitting function. Diffusion models (DMs) have also become popular choices for low-light image enhancement [104, 105, 106]. Diff-Retinex [105] formulates the low-light image enhancement problem into Retinex decomposition, and employs multi-path generative diffusion networks to reconstruct the normal-light Retinex probability distribution. A recent state-of-the-art approach presented in [106] decomposes images into high and low frequencies using wavelet transform. High frequencies are enhanced using a transformer-based pipeline, while the low frequencies undergo a diffusion process. This method achieves nearly 2.8dB improvement over the state-of-the-art transformer-based approach, e.g. LLFormer [100], and significantly better than INR-based method, NeRCo [108], on a real low-light image benchmarking dataset. The technique has been extended for video enhancement in [107]. The output of the enhancement typically depends on user preferences. This has been viewed as a one-to-many inverse problem, with attempts to solve it using Bayesian approaches. For example, a Bayesian Enhancement Model (BEM) [280] incorporates Bayesian Neural Networks (BNNs) to capture data uncertainty and produce diverse outputs. The method can be used with Transformers or Mamba as the architecture backbone.
@@ 3.3.4 Restoration
In our previous review paper [8], we categorized the work on restoration into several different types of distortions, including deblurring, denoising, dehazing, and mitigating atmospheric turbulence. Recent work, however, uses a unified network architecture to address these as inverse problems $y=hx+n$, where $x$ and $y$ are the ideal and observed data, respectively. $h$ is a degradation function, such as blur, and $n$ is additive noise. Often, the super-resolution task is also considered as an inverse problem, meaning $h$ includes the downsampling process. Note that although designed as a single network, the model is trained with each distorted dataset separately.
@@
The pioneering transformer-based method for image restoration, SwinIR [115], employs several concatenated Swin Transformer blocks [16]. SwinIR surpasses state-of-the-art CNN-based methods proposed up to the year 2021 in super-resolution and denoising tasks. The model is smaller and reconstructs fine details more effectively. Other two popular approaches that emerged in the same timeframe are Uformer [130] and Restormer [131]. Both incorporate Transformer blocks into hierarchical encoder-decoder networks, employing skip connections similar to those in U-Net. Their objective was to restore noisy images, sharpen blurry images, and remove rain. The networks focused on predicting the residual $R$ and obtaining the restored image $\hat{x}$ through $\hat{x}=y+R$. While their performance is very similar, Restormer has half the parameters of Uformer. More recently, GRL by Li et al. [119] exploits a hierarchy of features in a global, regional, and local range using different ways to compute self-attentions as an image often show similarity within itself in different scales and areas, which outperforms SwinIR and Restormer. Additionally, Fei et al. introduced the Generative Diffusion Prior [128] for unsupervised learning, aiming to model posterior distributions for image restoration and enhancement. VmambaIR [143] incorporates Mamba blocks into the U-Net architecture, achieving superior performance compared to SwinIR and Restormer in both visual quality and model size.
@@
For video restoration, the general framework comprises frame alignment, feature fusion and reconstruction. The process could be similar to image restoration but input multiple frames and run through the sequences in a sliding window manner to exploit the temporal information within a number of consecutive frames. Recently, Video Restoration Transformer (VRT) [103] and its improved version with a recurrent process (RVRT) [99], have emerged as the state of the arts for video super-resolution, deblurring, denoising, and frame interpolation. This method introduces temporal reciprocal self-attention in the transformer architecture and parallel warping using MLP. These innovations enable parallel computation and outperform the previous state-of-the-art methods by up to 2.16dB on benchmark datasets. FMA-Net [102] proposed multi-attention for joint video super-resolution and deblurring, achieving fast runtime with nearly 40% improvement over RVRT, and the restored quality was reported better by up to 3%.
@@
iii) Dehazing: Vision transformers for single image dehazing were proposed in DehazeFormer [135]. Similar to SUNet, it is a UNet-like architecture, but introduces Rescale Layer Normalization for better suit on improving contrast. The Fast Fourier Transform (FFT) has been employed in [140] due to the phase spectrum conveying more structural detail than the amplitude spectrum and demonstrating greater robustness to contrast distortion and noise. Then cross-attention between the RGB and YCbCr color spaces is applied. This approach achieves nearly 5 dB higher PSNR than DehazeFormer on a real-world smoke dataset. For video dehazing, Xu et al. [136] introduced a recurrent multi-range scene radiance recovery module with the space-time deformable attention. They also employ physics prior to inform haze attenuation. This method outperforms DehazeFormer by approximately 1dB.
iv) Mitigating atmospheric turbulence: Similar to dehazing, physics-inspired models have been widely developed to remove turbulence distortion [146, 148], while complex-valued CNNs have been proposed to exploit phase information [297]. There was also an attempt to use instance normalization (INR) to solve this problem, offering tile and blur correction [148]. However, diffusion models outperform on a single image [145], and transformer-based methods remain state-of-the-art for restoring videos [138, 139]. Mamba architecture, employed in [Hill2025MAMAT], outperforms Transformers and improves object detection performance. A recent review can be found in [299].